OpenStudy (anonymous):

The basketball team scored 12 points the first game and 47 points the eighth game of the season. How many points will they earn with this pattern on the 14th game of the season?

OpenStudy (goformit100):

@Wired

OpenStudy (anonymous):

In 8 games they increased the number of points they would get by 35 points. Take that difference and divide it by the 8 games that they have played. This will give you the number of points they increase by each game. Take that number and multiply it by the number of games between game 8 and 14th game (14-8=?). That will give you the total number of points you will have to add to 47 in order to find your final answer. Happy counting.

OpenStudy (anonymous):

f(x) = Ax+B f(1) = 12 = A(1)+B = A+B f(8) = 47 = A(8)+B = 8A+B Hint: From f(1) you can figure out that B=12-A Solve f(8) by substituting B, then you'll get A and the equation. Then solve for f(14).